OpenStudy (anonymous):

identify a slope for this line given points are: (0,-2) (3,-5) (6,-8), (9,-11)

OpenStudy (anonymous):

use all of this point to find slope (0,-2) (3,-5) (6,-8), (9,-11) the formula for slope m=(y2-y1)/(x2-x1) for ex using the first 2 points (0,-2) =(x2,y2),,(3,-5)=(x1,y1) m=(-2+5)/(0-3)=__? and for (3,-5), (6,-8) m=(-8+5)/(6-3)=___? and for (6,-8), (9,-11) m=(-8+11)/(6-9)=__? or m=(-11+8)/(9-6)=__?

OpenStudy (anonymous):

so the slope decrease by -12 as we change the points, is that right?

OpenStudy (anonymous):

hmm if thats the case yes, but i havent compute them yet i just show you haw to do the solution on finding the slope.... :D

OpenStudy (anonymous):

sorry, I was looking at the other problem

OpenStudy (anonymous):

So should I list all of the slopes between the points?

OpenStudy (anonymous):

Ok, I got it, Thanks for your help!

OpenStudy (anonymous):

i think you should graph them and find out whats the picture of it, i think its all equal slope

OpenStudy (anonymous):

ok good luck now, have fun :D

OpenStudy (anonymous):

So the slope is -1

OpenStudy (anonymous):

yes you got it.. :D